# Memorization Sinks: Isolating Memorization during LLM Training

### v1

@@ Abstract
Machine Learning, ICML
@@
1 Introduction
Large language models often memorize sequences that are repeated during pretraining (Carlini et al., 2023; Nasr et al., 2023), posing concerns in privacy, copyright, and membership inference. There is significant research on the problem of “unlearning” or removing memorized information from models post-hoc, including fine-tuning based approaches (Maini et al., 2022; Barbulescu & Triantafillou, 2024). However, presently all such methods present a substantial tradeoff between removing memorization and preserving general model capability.
We identify a core reason for the tradeoff: standard training induces mechanistic entanglement, where the same components support both generalization and memorization (Scherlis et al., 2025). This occurs when memorization relies on mechanisms also used for general language understanding. In a controlled setting, we show that repeated natural text sequences are memorized with such entanglement, making it challenging to remove them without harming general performance. We further prove that gradient descent has an implicit bias toward such entangled solutions, suggesting that mechanistic entanglement is inherent to current training methods.
We begin with a natural disentanglement attempt (Section 4; see also (Cloud et al., 2024)) by restricting gradient updates from repeated sequences to designated “memorized components”, while the remaining “general components” learn only from non-repeated data. This approach has two major flaws. First, it weakens generalization by depriving general components of all training signal from potentially high-quality repeated sequences. More subtly, generalization further degrades when memorized components are removed at inference: the model’s general capabilities co-adapt to the memorization neurons during training. Removing them post-hoc breaks this dependence and harms performance.
@@ 1 Introduction
The key insight lies in the difference in the training dynamics of generalizing and memorization signal, which has been studied in disparate lines of work. Generalizing signal is consistently amplified across training sequences (Chatterjee, 2020). However, memorization signals of different training sequences interfere with each other, resulting in a cyclical learning-forgetting dynamic (Toneva et al., 2018): learning when the repeated sequence is seen, but forgetting when training on a different sequence due to interference. In standard training, there is no separation of model components so this learning-forgetting dynamic occurs throughout the model leading to mechanistic entanglement. However, in MemSinks, this cycle is broken by allocating separate components per repeated sequence that are protected from interference with other sequences. As a result, memorized information is not reinforced in the shared components across repetitions of sequences.
How well does MemSinks disentangle memorization and generalization in practice? We train 360M and 1.7B SmolLM models (Allal et al., 2025) on the SlimPajama dataset (Shen et al., 2024), with repeated sequences drawn from TinyStories. Standard training leads to strong memorization where repeated sequences show significantly (i.e. 200x) lower loss than held-out ones. With MemSinks, dropping the memorization components closes over 50% of this loss gap, mitigating memorization. Furthermore, MemSinks (without memorization components) matches the validation loss of standard training and significantly outperforms a de-duplication baseline, preserving the benefits of repeated data for generalization. This provides a proof-of-concept that MemSinks can disentangle memorization from generalization in realistic settings (Section 5.2).
@@ 2 Related Works
With the discovery of memorization of sequences in large language models (Carlini et al., 2023; Nasr et al., 2023), there has been interest in techniques to remove memorization in LLMs post-hoc. One class of such methods involves further training of all language model parameters in order to reduce the likelihood of a memorized sequence. For example Thudi et al. (2022) presents the simple technique of simply training to increase the loss of memorized examples. Liu et al. (2022) further regularizes this by concurrently minimizing the loss on a retain set of validation set examples. Other works have examined using preference-based training to incentivize the replacement of a memorized sequence with a safe alternative (Zhang et al., 2024; Maini et al., 2023). However, currently all such unlearning methods are prone to degrading general model capabilities, beyond the desired unlearning target (Maini et al., 2023). Moreover, they require the computational expense of retraining all weights and in many cases access to additional data in the form of a retain set.
@@
A promising class of methods for removing memorization seek to localize specific model components, such as neurons, that are responsible for storing memorized sequences (Maini et al., 2023). Once such neurons are identified, they can simply be dropped out to ensure unlearning. These methods have the advantage of avoiding costly full model training and seek to avoid model degradation by minimizing the number of parameters changed. Localization techniques have achieved success in removing personally identifying information (PII) (Chen et al., 2024), as well as when there exist distinctive memorization triggers (Stoehr et al., 2024). However, all such methods present a tradeoff between removing memorization and preserving model performance (Chang et al., 2024b).
@@
Sparse mixture-of-experts models (S-MOE) have been studied in prior works as a means to expand model capacity without increasing compute costs (Shazeer et al., 2017). Like MemSinks, MOE models selectively activate fully-connected neurons depending on the input token. However, the selection of neurons in SMOE is typically learned subject to load-balancing constraints (as opposed to explicitly enforced as in MemSinks). Consequently, it is unclear what degree of localization is achieved by SMOE. Zoph et al. (2022) find that experts generally specialize based off of low-level syntactic structures rather than semantic features. Dai et al. (2024) find that there is often redundancy across experts in MOEs, challenging the idea that they reliably localize information within experts. Other works (Gururangan et al., 2021; Park et al., 2025) provide some evidence of localization in MOE models: however they measure localization relative to coarse attributes such as language and general topics. In this work we study the far more fine-grained problem of localizing memorization of specific documents or sequences.
@@ 3.2 Empirical Observations
In Figure 2(c), we plot the validation and repeated sequence losses of a model trained on TS-Repetition. We see that the loss on repeated sequences and the validation set descend simultaneously. The simultaneous learning of memorization and generalization seen here underscores the tight interplay between natural sequence memorization and general capabilities, contrasting from settings such as label noise where memorization occupies a distinct phase of training (Li et al., 2020).
@@ 4.2 Empirical Findings
We observe that the performance of gradient masking is inferior to a standard model, even before memorization neurons are removed (Figure 3). This observation renders gradient masking impractical, as it significantly worsens the model’s general capabilities. Our findings here highlight that some “shared” neurons – those that are updated by all sequences – are essential to maximally aggregate generalizing features. However, this suggests a potential tradeoff between generalization and disentangling memorization: it is unclear how to ensure that shared neurons do not implement memorization as well.
@@ 5 Channeling Training Dynamics with MemSinks
To address this, we leverage the difference in training dynamics of memorization and generalization. During training, generalizing signals are reinforced across distinct sequences and are broadly amplified (Chatterjee, 2020). Sequence-specific memorization, on the other hand, experiences interference from other examples and is gradually forgotten. As a result, repeated sequences are cyclically learned and forgotten throughout the training process, thereby becoming defused throughout the model.
@@
We implement MemSinks in the hidden layer of transformer MLPs, as prior works have found MLPs play a crucial role in storing memorization (Nanda et al., 2023; Geva et al., 2021). We set a fraction $g$ of the hidden layer neurons in each layer as generalization neurons (which are activated across all sequences), while the remaining $1-g$ fraction serve as memorization sink neurons. We assign each sequence in the pretraining dataset an ID and mask the memorization sink neurons deterministically as a function of the sequence ID. During all evaluations, we remove memorization sink neurons. Further implementation details are provided in Appendix G.
@@ Appendix A Extended Related Works
A large body of work has focused on unlearning or forgetting memorized information from neural models, especially in the classification domain. This includes methods such as SISA (Bourtoule et al., 2021) that exactly unlearn information by maintaining multiple model copies, and a recent influx of approximate unlearning approaches (Triantafillou et al., 2023) that aim to perform post-hoc procedures on a model in order to remove information in question. The study of memorization in classification settings is connected to label noise and atypical examples (Harutyunyan et al., 2020; Feldman & Zhang, 2020). On the other hand, in this work, we concentrate on more natural or typical memorized points which often arise in the large language model setting.
@@ Appendix B Implementation Details of TinyStories Training
Min Learning Rate Max Learning $\frac{\text{Max Learning Rate}}{10}$
@@ Appendix C Implementation Details of Post-hoc Localization Techniques
We generally follow the methodology used in (Chang et al., 2024b) and directly used their code as released online. We restrict our attention to their Hard-Concrete and Integrated Gradients methods presented in the papers.
@@
We tuned $\lambda$, the $\ell_{1}$ loss coefficient used in training the pruning mask $M$ over the values $\{100,500,1000\}$ on a tuning set of 5 sequences. Additionally, we tuned the number of pruning iterations in the range $\{1000,2000,4000\}$. The remainder of hyperparameters were set to the optimal values reported by (Chang et al., 2024a). We tuned relative to the lowest validation loss achieved after dropping out the identified neurons.
@@
Following the computation of mask scores by either Hard Concrete or attribution scores by Integrated Gradients, we sorted the neurons in each layer by these scores. Given a dropout parameter $r$, we dropped out an $r$ proportion of the neurons in each layer, as was performed in (Chang et al., 2024a).
@@ Appendix D Disentanglement in Standard Training
We will consider a dataset with $n$ examples seen once $\mathcal{D}_{\textrm{once}}=\{(s_{i},y_{i})\}_{i=1}^{N}$ and a memorization sequence $(s^{\textrm{mem}},y^{\textrm{mem}})$ that is repeated $k$ times. Thus, explicitly, our training dataset consists of $\mathcal{D}_{\textrm{pre}}=\mathcal{D}_{\textrm{once}}\cup(\cup_{k}(s^{\textrm{mem}},y^{\textrm{mem}}))$. For the purposes of this theory, we will assume that the $\mathbb{R}^{d_{\textrm{emb}}}$ can be partitioned into a semantic subspace $S_{\textrm{semantic}}$ and a memorization subspace $S_{\textrm{noise}}$; explicitly we will write that $\mathbf{\phi(s^{i})}=\begin{bmatrix}\mathbf{\phi(s)}^{(i)}_{\textrm{mem}}&\mathbf{\phi(s)}^{(i)}_{\textrm{mem}}\end{bmatrix}^{\top}$. We will make the assumption that $\forall i\in[1,N]$ $(\mathbf{\phi(s)}_{\textrm{mem}}^{(i)})^{\top}\mathbf{\phi(s)}_{\textrm{mem}}^{(\textrm{mem})}=0$. This can be interpreted as enabling us to uniquely identify a particular example in the noise subspace which would facilitate disentanglement of memorization. For convenience, we will consider the matrix implemented by $\mathbf{W}=\mathbf{W_{\text{proj}}}\mathbf{W_{\text{fc}}}$ and we will denote $\mathbf{W}^{*}=\textrm{argmin}_{L(\mathbf{W},\mathcal{D_{\textrm{once}}})=0}\left\lVert{W}\right\rVert_{F}$ (i.e. the minimum norm solution when $\mathbf{W}^{*}$ when the memorization example is not seen). We will assume that $\textrm{rank}(\mathbf{W^{*}})>2$. We will assume that we can achieve 00 loss using only the subspace $S_{\textrm{semantic}}$. In particular, denoting the singular value decomposition of $\mathbf{W}^{*}=\mathbf{U}^{*}\Sigma(\mathbf{V^{*}})^{\top}$, this translates to the condition that $\textrm{span}(\mathbf{V}^{*})\subseteq S_{semantic}$. We will denote the term $\Delta_{\textrm{mem}}=y^{\textrm{mem}}-f(\mathbf{\phi(s^{\textrm{mem}})})$ (the change in prediction necessary to memorize $s^{\textrm{mem}}$) and assume $||\Delta_{\textrm{mem}}||>1$. We will assume that all the singular values of $\sigma_{i}(\mathbf{W^{*}})<\frac{1}{2k}$ $\forall i\in[1,\textrm{rank}(\mathbf{W^{*}})]$ (i.e. the generalizing solution has low norm) and that the memorization activations are unit norm $||(\mathbf{\phi(\mathbf{s})^{\textrm{mem}}_{\textrm{mem}}}||=1$). Finally, we will assume that the $\Delta_{\textrm{mem}}\perp\text{col}(\mathbf{W^{*}})$.
@@ D.1 Analysis of Natural Sequence Memorization
A learned parameter $\mathbf{\hat{W}_{\textrm{dis}}}$ implements the Disentangled Memorizing Solution if $\mathbf{\hat{W}_{\textrm{dis}}}$ can be written as $\mathbf{W^{*}}+c\mathbf{u}\mathbf{v}^{\top}$ where $\mathbf{v}\in S_{\textrm{sem}}$ and $\mathbf{\hat{W}_{\textrm{dis}}}$ achieves 00 training loss on $\mathcal{D}_{\textrm{pre}}$
@@
A learned parameter $\mathbf{\hat{W}_{\textrm{ent}}}$ implements the entangled memorizing solution if $\mathbf{\hat{W}_{\textrm{ent}}}$ achieves 0 training loss on $\mathcal{D}_{\textrm{pre}}$ and $\mathbf{\hat{W}_{\textrm{ent}}}$ can be written as $\sum\limits_{i=1}^{k}\sigma^{(\Delta)}_{i}\tilde{\mathbf{u}}_{i}\mathbf{v_{i}}^{\top}+\sigma_{k+1}^{\Delta}\mathbf{\tilde{u}_{k+1}}\mathbf{v}_{k+1}$ where the $\mathbf{v}_{1},...\mathbf{v}_{i}$ are the original right singular vectors of $\mathbf{W^{*}}$ and $\mathbf{\tilde{u}}_{i}$ are a potentially shifted set of singular vectors and $\mathbf{\tilde{u}}_{i+1},\mathbf{\tilde{v}_{i+1}}$ are a new set of vectors and $\forall i\in[1,k]$ we have that $\sigma_{i}^{\Delta}\leq\sigma(\mathbf{W}^{*})_{i}+\frac{\left\lVert\Delta_{\textrm{mem}}\right\rVert}{k+1}$ and $\sigma_{k+1}^{\Delta}\leq\frac{||\Delta_{\textrm{mem}}||}{k+1}$ where $k=\textrm{rank}(\mathbf{W^{*}})$
Intuitively, entangled memorizing solution reuses and shifts vectors that are used by the generalizing solution. We impose the condition on the singular values to model that the implementation of memorization is “diffused” across multiple of the generalizing semantically meaningful features.
@@
    The effective linear predictor converges to a min-norm solution, formally $\lim_{t\rightarrow\infty}\mathcal{\beta}(t)=\beta^{*}$ where $\beta^{*}=\textrm{argmin}_{\beta\in I(D_{\textrm{pre}})}||\beta||_{F}$
Using the result from Theorem D.5, the proof of Theorem D.3 is straightforward and simply requires comparing the Frobenius norms of the disentangled and entangled memorizing solutions. In particular, because $||\phi(\mathbf{s}^{(i)\top})||_{2}=1$, we must have that $||\mathbf{W_{\textrm{dis}}}||_{F}^{2}=||\mathbf{W^{*}}||_{F}^{2}+||\left\lVert\Delta_{\textrm{mem}}\right\rVert_{2}^{2}$ (the remaining singular value comes from the orthogonal rank-1 term $\mathbf{u}\mathbf{v}^{\top}$). On the other hand, consider the entangled memorizing solution. We have that $||\mathbf{W_{\textrm{ent}}}||_{F}^{2}\leq||\mathbf{W}^{*}||_{F}^{2}+2\sum\limits_{i=1}^{k}\sigma(\mathbf{W^{*}})_{i}\frac{||\Delta_{\textrm{mem}}||}{k+1}+\frac{||\Delta_{\textrm{mem}}||^{2}}{k+1}\leq||\mathbf{W^{*}}||_{F}^{2}+2\frac{||\Delta_{\textrm{mem}}||_{2}^{2}}{k}\leq||\mathbf{W^{*}}||_{F}^{2}+||\Delta_{\textrm{mem}}||_{2}^{2}$

### v2

@@ Abstract
1  Introduction
Large language models often memorize sequences that are repeated during pretraining (Carlini et al., 2023; Nasr et al., 2023), posing concerns in privacy, copyright, and membership inference. There is significant research on the problem of “unlearning” or removing memorized information from models post-hoc, including fine-tuning based approaches Maini et al. (2022); Barbulescu & Triantafillou (2024). However, presently all such methods present a substantial tradeoff between removing memorization and preserving general model capability. We identify a core reason for the tradeoff: standard training induces mechanistic entanglement, where the same components support both generalization and memorization Scherlis et al. (2025). This occurs when memorization relies on mechanisms also used for general language understanding. In a controlled setting, we show that repeated natural text sequences are memorized with such entanglement, making it challenging to remove them without harming general performance. We further prove that gradient descent has an implicit bias toward such entangled solutions, suggesting that mechanistic entanglement is inherent to current training methods.
We begin with a natural disentanglement attempt (Section 4; see also Cloud et al. (2024)) by restricting gradient updates from repeated sequences to designated “memorized components”, while the remaining “general components” learn only from non-repeated data. This approach has two major flaws. First, it weakens generalization by depriving general components of all training signal from potentially high-quality repeated sequences. More subtly, generalization further degrades when memorized components are removed at inference: the model’s general capabilities co-adapt to the memorization neurons during training. Removing them post-hoc breaks this dependence and harms performance.
@@ 1  Introduction
The key insight lies in the difference in the training dynamics of generalizing and memorization signal, which has been studied in disparate lines of work. Generalizing signal is consistently amplified across training sequences Chatterjee (2020). However, memorization signals of different training sequences interfere with each other, resulting in a cyclical learning-forgetting dynamic (Toneva et al., 2018): learning when the repeated sequence is seen, but forgetting when training on a different sequence due to interference. In standard training, there is no separation of model components so this learning-forgetting dynamic occurs throughout the model leading to mechanistic entanglement. However, in MemSinks, this cycle is broken by allocating separate components per repeated sequence that are protected from interference with other sequences. As a result, memorized information is not reinforced in the shared components across repetitions of sequences.
How well does MemSinks disentangle memorization and generalization in practice? We train 360M and 1.7B SmolLM models Allal et al. (2025) on the SlimPajama dataset Shen et al. (2024), with repeated sequences drawn from TinyStories. Standard training leads to strong memorization where repeated sequences show significantly (i.e. 200x) lower loss than held-out ones. With MemSinks, dropping the memorization components closes over 50% of this loss gap, mitigating memorization. Furthermore, MemSinks (without memorization components) matches the validation loss of standard training and significantly outperforms a de-duplication baseline, preserving the benefits of repeated data for generalization. This provides a proof-of-concept that MemSinks can disentangle memorization from generalization in realistic settings (Section 5.2).
@@ 2  Related Works
With the discovery of memorization of sequences in large language models Carlini et al. (2023); Nasr et al. (2023), there has been interest in techniques to remove memorization in LLMs post-hoc. One class of such methods involves further training of all language model parameters in order to reduce the likelihood of a memorized sequence. For example Thudi et al. (2022) presents the simple technique of simply training to increase the loss of memorized examples. Liu et al. (2022) further regularizes this by concurrently minimizing the loss on a retain set of validation set examples. Other works have examined using preference-based training to incentivize the replacement of a memorized sequence with a safe alternative Zhang et al. (2024); Maini et al. (2023). However, currently all such unlearning methods are prone to degrading general model capabilities, beyond the desired unlearning target Maini et al. (2023). Moreover, they require the computational expense of retraining all weights and in many cases access to additional data in the form of a retain set.
@@
A promising class of methods for removing memorization seek to localize specific model components, such as neurons, that are responsible for storing memorized sequences Maini et al. (2023). Once such neurons are identified, they can simply be dropped out to ensure unlearning. These methods have the advantage of avoiding costly full model training and seek to avoid model degradation by minimizing the number of parameters changed. Localization techniques have achieved success in removing personally identifying information (PII) Chen et al. (2024), as well as when there exist distinctive memorization triggers Stoehr et al. (2024). However, all such methods present a tradeoff between removing memorization and preserving model performance Chang et al. (2024b).
@@
Sparse mixture-of-experts models (S-MOE) have been studied in prior works as a means to expand model capacity without increasing compute costs Shazeer et al. (2017). Like MemSinks, MOE models selectively activate fully-connected neurons depending on the input token. However, the selection of neurons in SMOE is typically learned subject to load-balancing constraints (as opposed to explicitly enforced as in MemSinks). Consequently, it is unclear what degree of localization is achieved by SMOE. Zoph et al. (2022) find that experts generally specialize based off of low-level syntactic structures rather than semantic features. Dai et al. (2024) find that there is often redundancy across experts in MOEs, challenging the idea that they reliably localize information within experts. Other works Gururangan et al. (2021); Park et al. (2025) provide some evidence of localization in MOE models: however they measure localization relative to coarse attributes such as language and general topics. In this work we study the far more fine-grained problem of localizing memorization of specific documents or sequences.
@@ 3.2 Empirical Observations
In Figure 2(c), we plot the validation and repeated sequence losses of a model trained on TS-Repetition. We see that the loss on repeated sequences and the validation set descend simultaneously. The simultaneous learning of memorization and generalization seen here underscores the tight interplay between natural sequence memorization and general capabilities, contrasting from settings such as label noise where memorization occupies a distinct phase of training Li et al. (2020).
@@ 4.2 Empirical Findings
We observe that the performance of gradient masking is inferior to a standard model, even before memorization neurons are removed (Figure 3). This observation renders gradient masking impractical, as it significantly worsens the model’s general capabilities. Our findings here highlight that some “shared" neurons – those that are updated by all sequences – are essential to maximally aggregate generalizing features. However, this suggests a potential tradeoff between generalization and disentangling memorization: it is unclear how to ensure that shared neurons do not implement memorization as well.
@@ 5  Channeling Training Dynamics with MemSinks
To address this, we leverage the difference in training dynamics of memorization and generalization. During training, generalizing signals are reinforced across distinct sequences and are broadly amplified Chatterjee (2020). Sequence-specific memorization, on the other hand, experiences interference from other examples and is gradually forgotten. As a result, repeated sequences are cyclically learned and forgotten throughout the training process, thereby becoming defused throughout the model.
@@
We implement MemSinks in the hidden layer of transformer MLPs, as prior works have found MLPs play a crucial role in storing memorization Nanda et al. (2023); Geva et al. (2021). We set a fraction $g$ of the hidden layer neurons in each layer as generalization neurons (which are activated across all sequences), while the remaining 1g1-g fraction serve as memorization sink neurons. We assign each sequence in the pretraining dataset an ID and mask the memorization sink neurons deterministically as a function of the sequence ID. During all evaluations, we remove memorization sink neurons. Further implementation details are provided in Appendix G.
@@ Appendix A Extended Related Works
A large body of work has focused on unlearning or forgetting memorized information from neural models, especially in the classification domain. This includes methods such as SISA (Bourtoule et al., 2021) that exactly unlearn information by maintaining multiple model copies, and a recent influx of approximate unlearning approaches (Triantafillou et al., 2023) that aim to perform post-hoc procedures on a model in order to remove information in question. The study of memorization in classification settings is connected to label noise and atypical examples Harutyunyan et al. (2020); Feldman & Zhang (2020). On the other hand, in this work, we concentrate on more natural or typical memorized points which often arise in the large language model setting.
@@ Appendix B Implementation Details of TinyStories Training
Min Learning Rate $\frac{\text{Max Learning Rate}}{10}$
@@ Appendix C Implementation Details of Post-hoc Localization Techniques
We generally follow the methodology used in Chang et al. (2024b) and directly used their code as released online. We restrict our attention to their Hard-Concrete and Integrated Gradients methods presented in the papers.
@@
We tuned $\lambda$, the $\ell_{1}$ loss coefficient used in training the pruning mask $M$ over the values $\{100,500,1000\}$ on a tuning set of 5 sequences. Additionally, we tuned the number of pruning iterations in the range $\{1000,2000,4000\}$. The remainder of hyperparameters were set to the optimal values reported by Chang et al. (2024a). We tuned relative to the lowest validation loss achieved after dropping out the identified neurons.
@@
Following the computation of mask scores by either Hard Concrete or attribution scores by Integrated Gradients, we sorted the neurons in each layer by these scores. Given a dropout parameter $r$, we dropped out an $r$ proportion of the neurons in each layer, as was performed in Chang et al. (2024a).
@@ Appendix D Disentanglement in Standard Training
We will consider a dataset with $n$ examples seen once $\mathcal{D}_{\textrm{once}}=\{(s_{i},y_{i})\}_{i=1}^{N}$ and a memorization sequence $(s^{\textrm{mem}},y^{\textrm{mem}})$ that is repeated $k$ times.Thus, explicitly, our training dataset consists of $\mathcal{D}_{\textrm{pre}}=\mathcal{D}_{\textrm{once}}\cup(\cup_{k}(s^{\textrm{mem}},y^{\textrm{mem}}))$. For the purposes of this theory, we will assume that the $\mathbb{R}^{d_{\textrm{emb}}}$ can be partitioned into a semantic subspace $S_{\textrm{semantic}}$ and a memorization subspace $S_{\textrm{noise}}$; explicitly we will write that $\mathbf{\phi(s^{i})}=\begin{bmatrix}\mathbf{\phi(s)}^{(i)}_{\textrm{mem}}&\mathbf{\phi(s)}^{(i)}_{\textrm{mem}}\end{bmatrix}^{\top}$. We will make the assumption that $\forall i\in[1,N]$ $(\mathbf{\phi(s)}_{\textrm{mem}}^{(i)})^{\top}\mathbf{\phi(s)}_{\textrm{mem}}^{(\textrm{mem})}=0$. This can be interpreted as enabling us to uniquely identify a particular example in the noise subspace which would facilitate disentanglement of memorization. For convenience, we will consider the matrix implemented by $\mathbf{W}=\mathbf{W_{\text{proj}}}\mathbf{W_{\text{fc}}}$ and we will denote $\mathbf{W}^{*}=\textrm{argmin}_{L(\mathbf{W},\mathcal{D_{\textrm{once}}})=0}\left\lVert{W}\right\rVert_{F}$ (i.e. the minimum norm solution when $\mathbf{W}^{*}$ when the memorization example is not seen). We will assume that $\textrm{rank}(\mathbf{W^{*}})>2$. We will assume that we can achieve 0 loss using only the subspace $S_{\textrm{semantic}}$. In particular, denoting the singular value decomposition of $\mathbf{W}^{*}=\mathbf{U}^{*}\Sigma(\mathbf{V^{*}})^{\top}$, this translates to the condition that $\textrm{span}(\mathbf{V}^{*})\subseteq S_{semantic}$. We will denote the term $\Delta_{\textrm{mem}}=y^{\textrm{mem}}-f(\mathbf{\phi(s^{\textrm{mem}})})$ (the change in prediction necessary to memorize $s^{\textrm{mem}}$) and assume $\Delta_{\textrm{mem}}||>1$. We will assume that all the singular values of $\sigma_{i}(\mathbf{W^{*}})<\frac{1}{2k}$ $\forall i\in[1,\textrm{rank}(\mathbf{W^{*}})]$ (i.e. the generalizing solution has low norm) and that the memorization activations are unit norm $||(\mathbf{\phi(\mathbf{s})^{\textrm{mem}}_{\textrm{mem}}}||=1$). Finally, we will assume that the $\Delta_{\textrm{mem}}\perp\text{col}(\mathbf{W^{*}})$.
@@ D.1 Analysis of Natural Sequence Memorization
A learned parameter $\mathbf{\hat{W}_{\textrm{dis}}}$ implements the Disentangled Memorizing Solution if $\mathbf{\hat{W}_{\textrm{dis}}}$ can be written as $\mathbf{W^{*}}+c\mathbf{u}\mathbf{v}^{\top}$ where $\mathbf{v}\in S_{\textrm{sem}}$ and $\mathbf{\hat{W}_{\textrm{dis}}}$ achieves 0 training loss on $\mathcal{D}_{\textrm{pre}}$
@@
[…]
Intuitively, entangled memorizing solution reuses and shifts vectors that are used by the generalizing solution. We impose the condition on the singular values to model that the implementation of memorization is “diffused" across multiple of the generalizing semantically meaningful features.
@@
[…]